Grade 9Magical Mathematics[Interesting Approach]

Let a+b+c+ab+bc+ca+abc = 1000

Then The Possible Value Of a+b+c Is

A) 28

B) 36

C) 39

D) 42

PRAKHAR  GUPTA
13 Years agoGrade 9

1 Answer

jitender lakhanpal
13 Years ago

hi prakhar

the expression is (a+1)(b+1)(c+1) - 1 = 1000

(a+1)(b+1)(c+1)=1001 = 7*11*13

from here we can conclude a = 6, b= 10 , c= 12

can be the possible values

We are  here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

jitender ''

askiitians expert